OpenStudy (anonymous):

NEED HELP WILL FAN AND MEDAL FOR FIRST ANSWER. CORRECT ASNWER = 2 MEDALS AND 2 FANS. WILL GET MY SISTERS ON TO FAN AND MEDAL YOU. ANSWER FAST, THOUGH!

OpenStudy (anonymous):

Here is question incase you cannot see the picture.

OpenStudy (anonymous):

@ginger1427 What is the solution to the system of equations? 2x + y = -3 x - 2y = -4

OpenStudy (anonymous):

@Smart_Owl_Student ^^

OpenStudy (anonymous):

What there asking is 2 x what + what + -3

OpenStudy (anonymous):

thats for number 1

OpenStudy (anonymous):

@Smart_Owl_Student I'm confused ?

OpenStudy (anonymous):

number 2: what - 2 x what = -4

OpenStudy (anonymous):

just study that... it will come to u :-)

OpenStudy (anonymous):

@Smart_Owl_Student I'm still confused .. What are you saying?

OpenStudy (anonymous):

@Smart_Owl_Student I have solved for x in the first equation - x = -1/2y - 3/2 and plugged x back in to solve for y and I get -1y - 3 + y = -3 which cancels everything out. I know I am doing something wrong but can't figure out

OpenStudy (anonymous):

@Smart_Owl_Student it still hasn't come to me...help..?